- Status Report
- August 10, 2022
SOHO Daily Meeting Minutes for Tuesday, 7 May 2002
* RHESSI target for today is AR 9934; target for tomorrow
is AR 9934.
* JOP104 target for today is AR 9919, 9926 (and 9930);
target for tomorrow is the same.
* PMI has been scheduled to perform corrective maintenance on
PDU-D16 on May 7 from approx. 10:30 am until 2:30 pm. This unit
is one of two PDU’s that feed the ISTP areas in building 3 rooms
C130, C150, C156 and C160. All equipment on this PDU must be
shut down and powered off prior to the maintenance start time.
THERE WILL BE NO COMMANDING OR TELEMETRY IN THE EOF DURING THIS
TIME. ALL DATA WILL BE RECOVERABLE.
Spacecraft Status: Nominal
126/1933 VIRGO sensors SPM-A and SPM-B latched-up for 30
minutes. Impacts TBD.
Accomplished Activities: VIRGO
Planned Activities: VIRGO, CELIAS, RSL
Upcoming Operations: None
126/1605 Low SPE value received during D27 support using
RCVR 2. Problem cleared following a second carrier
sweep. Commanding activities delayed 23 minutes.
126/1607 Lost 16 minutes 6 seconds of recoverable TLM, and
14 minutes of potential MDI high rate data, due to
a SCP failure at Goldstone.
126/2207 Command corrupted bewtween D27 antenna and the
spacecraft. Commanding activities (includig NRT)
delayed 6 minutes 15 seconds.
127/0847 CMA failure resulted in commanding activities being
delayed 14 minutes.
05/06 DOY 126 – 05/07 DOY 127
EIT daily On-Line Solar Status and observations:
LASCO daily On-Line Solar Status and observations:
Web Page for Planning:
CDS: Nominal. JOP104, sunspot oscillation study on AR9934,
engineering. Tomorrow: JOP136, engineering, and either
sunspot oscillations study or active region maps.
SUMER: Nominal. JOP104 continues.
UVCS: Nominal. Coronal hole observations and mini-synoptic.
Tomorrow: coronal hole observations and synoptic.
LASCO: Nominal. C2/C3 synoptics.
EIT: Nominal. Half-resolution 195 CME watch and synoptic.
MDI: Nominal. Full disk magnetograms and dopplergrams for continuous
TRACE: Nominal. Observing the RHESSI target in 195A.